# Differentially private and decentralized randomized power method

### v2

@@ 5.1 Decentralized and private power method using distributed DP:
We introduce a federated version of the Privacy-Preserving Power Method in Algorithm 2. This version reduces the noise variance by a factor of $O(sp\log d\,\|\bm{U}\|_{\infty}^{2})$ compared to the method in Balcan et al. (2016). The improvement comes from the fact that we emulate centralized Differential Privacy (DP) over secure channels, avoiding the higher noise required in local DP settings with public channels, and that we use our generally tighter sensitivity bound.
We demonstrate the equivalence between the decentralized Algorithm 2 and the centralized Algorithm 1 in Theorem 5.2 (proof provided in Appendix A.5). We also give a simplified definition of the Secure Aggregation algorithm.
@@
We denote by $SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1{\leq}i{\leq}s\})$ the Secure Aggregation of locally held matrices $\bm{Y}_{\ell}^{(i)}$ over the set of users represented by indices $\{i|1{\leq}i{\leq}s\})$. It is equivalent to computing $\bm{Y}_{\ell}=\sum_{i=1}^{s}\bm{Y}_{\ell}^{(i)}$ over a secure channel.
  Input: distributed matrices $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, private parameters $\epsilon,\delta$.
  Output: approximated eigen-space $\bm{X}^{L}\in\mathbb{R}^{n\times p}$, with orthonormal columns.
  Initialization: orthonormal $\bm{X}^{0}\in\mathbb{R}^{n\times p}$ by QR decomposition on a random Gaussian matrix $\bm{G}_{0}$; noise variance parameter $\nu=\epsilon^{-1}\sqrt{\frac{4L\log(1/\delta)}{s}}$;
  for $\ell=1$ to $L$ do
      1. The central node broadcasts $\bm{X}^{\ell-1}$ to all $s$ computing nodes;2. Computing node $i$ computes $\bm{Y}^{(i)}_{\ell}=\bm{A}^{(i)}\bm{X}^{\ell-1}+\bm{G}^{(i)}_{\ell}$ with $\bm{G}^{(i)}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\nu^{2})^{n\times p}$;3. The central node computes with the clients $\bm{Y}_{\ell}=SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1\leq i\leq s\})$. 4. The central node computes QR factorization $\bm{Y}_{\ell}=\bm{X}^{\ell}\bm{R}_{\ell}$.
  end for
@@ 5.2 Communication and computation cost analysis:
We denote by $s$ the number of shards, such that $\sum_{i=1}^{s}\bm{A}^{(i)}=\bm{A}\in\mathbb{R}^{n\times n}$. Table 1 presents the communication and computations cost overhead introduced by the polylogarithmic version of Secure Aggregation that we employ (Bell et al., 2020).
@@ 6 Empirical comparison of the proposed bounds
We introduced both runtime-dependent and runtime-independent convergence bounds for our algorithm. The runtime-dependent bound depends on the matrix $\bm{A}$ and the iteratively computed matrices $\bm{X}_{l}$, whereas the runtime-independent bound (Theorem 4.1) only depends on $\bm{A}$. To illustrate the practical impact of our changes, we focus on an application in recommender systems and compare how our algorithm perform in this context. We will then use a statistical approximation to see how the runtime-dependent bounds compare at the first step of the algorithm, regardless of the application.
@@ 6.1 Application to recommender systems
State-of-the-art recommender systems like GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) utilize singular value decomposition (SVD) or eigenvalue decomposition as part of their algorithms. Specifically, they represent the dataset (comprised of user-item interactions) using a normalized adjacency matrix $\tilde{\bm{R}}$, from which they compute a normalized item-item matrix $\tilde{\bm{P}}$. To get rid of noise in $\tilde{\bm{R}}$, they compute an ideal low-pass filter based on the top-$p$ eigenvectors of $\tilde{\bm{P}}$ and apply it to $\tilde{\bm{R}}$. Detailed definitions are provided in Appendix A.7.
Table 2 presents key statistics of popular recommendation datasets used in the literature. Depending on the use case, these datasets may be held by a central curator, or distributed among the users of a recommender system. Accordingly, either the centralized PPM (Algorithm 1) or the decentralized PPM (Algorithm 2) can be applied to compute the leading eigenvectors. To demonstrate the practicality of our proposed methods, we focus on a decentralized setting where each user has access to part of the dataset and collaborate with other users to compute the desired eigenvectors.
@@
Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{prior}$ (existing) ${\hat{\Delta}_{l}}$ (ours)
Amazon-book 52,643 91,599 2,984,108 0.33 52.95 1.12 ×$\times\sqrt{p}$ 1
MovieLens 71,567 10,677 7,972,582 0.39 32.84 1.19 ×$\times\sqrt{p}$ 1
EachMovie 74,425 1,649 2,216,887 0.49 21.89 1.33 ×$\times\sqrt{p}$ 1
Jester 54,906 151 1,450,010 0.73 7.78 1.64 ×$\times\sqrt{p}$ 1
6.1.1 Empirical comparison of the runtime-independent bounds
As discussed in Section 4.1, our proposed runtime-independent bound is tighter than those proposed in Hardt & Price (2014); Balcan et al. (2016) when $\min(1,\mu_{1}(\bm{A}))\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({d})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$. Table 2 shows that this condition holds for popular recommender system datasets. As we can deduce from Table 2, the proposed bound theoretically allows us to converge to solutions with much smaller $\eta$ values, especially when the desired number of factors $p$ is large.
6.1.2 Practical utility of the proposed algorithms
We saw that the proposed runtime-independent bounds were practically tighter than the previous one for the task of interest. We demonstrate that our decentralized PPM can privately compute the top-$p$ eigenvectors $\bm{U}_{p}$ of the user-normalized item-item matrix $\tilde{\bm{P}}$ (proof in Appendix A.1):
@@ 6.1.2 Practical utility of the proposed algorithms
Algorithm 2 with $\Delta_{l}=\sqrt{2}\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ can compute $\bm{U}_{p}$, the top-$p$ eigenvectors of $\tilde{\bm{P}}$ in a decentralized setting under $(\epsilon,\delta)$-Differential Privacy.
Figure 1: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the EachMovie dataset.
@@
Both GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) use $\bm{U}_{p}$ to compute the ideal low pass filter and apply it to the interaction matrix $\bm{R}$, giving a filtered interaction matrix $\bm{R}_{p}$. To illustrate the practicality of our proposed method, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using either our proposed decentralized PPM, or the PPM versions proposed by  Balcan et al. (2016). We then use $\tilde{\bm{U}}_{p}$ to compute an approximation of $\bm{R}_{p}$ (noted $\tilde{\bm{R}_{p}}$), and compute the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to each version.
Figure 1 illustrates the impact of the privacy parameter $\epsilon$ on the approximation error for the EachMovie dataset, with $p=32$ and $L=3$. Our decentralized PPM achieves a relatively low approximation error (approximately 1/10 of the random guessing error) for practical values of $\epsilon$ (5-10). In contrast, the methods from Balcan et al. (2016) require $\epsilon$ values of at least $10^{3}$ to achieve comparable errors, which does not provide meaningful privacy protection. Additional comparisons and details are provided in Appendix A.7.1.
6.2 Matrix-agnostic comparison of the runtime-dependent bounds
At iteration $l=0$, the leading singular vectors are initialized as $\bm{X}^{0}=\text{Q}(\bm{\Omega})$, where $\bm{\Omega}\sim\mathcal{N}(0,1)^{n\times k}$ and $\text{Q}(\bm{\Omega})$ denotes the orthonormal matrix obtained from the QR decomposition of $\bm{\Omega}$. Since $\bm{X}^{0}$ is independent of the matrix $\bm{A}$, it allows us to assess the relative tightness of the proposed bound $\hat{\Delta}_{0}$ compared to the previous bound $\Delta_{0}^{prior}$ at the first step.
Let $r(k,n)$ be the ratio of the expected values of the two bounds, as a function of the desired number of eigenvectors $k$ and the matrix dimension $n$:
$r(k,n)=\frac{\mathbb{E}[\Delta_{0}^{prior}]}{\mathbb{E}[\hat{\Delta}_{0}]}.$ (18)
6.2.1 Theoretical approximation of $r(k,n)$:
We derive an asymptotic approximation of $r(k,n)$ using Approximation 6.2 (detailed in Appendix A.4).
Approximation 6.2.
[…]
$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(kn)\frac{k^{2}}{n^{2}}}+\frac{k\mu}{n}\right)$ (19)
$\mathbb{E}[\hat{\Delta}_{l}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(n)\frac{k}{n^{2}}}+\frac{k\mu}{n}\right).$ (20)
6.2.2 Empirical approximation of $r(k,n)$:
Figure 2: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=8000$ and $k$ ranging between 64 and 4000, with a step of 64. Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm. The theoretical ratios are estimated using Theorem 6.2.
To compare the tightness of our proposed bound $\hat{\Delta}_{0}$ to the previous bound $\Delta_{0}^{prior}$, we perform both theoretical and empirical analyses at the first iteration, where the noise scaling depends only on the random initialization $\bm{X}^{0}$. Since $\bm{X}^{0}$ is independent of $\bm{A}$, we can estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$ by sampling random Gaussian matrices $\bm{X}^{0}$.
Experiment: $r(k,n)$ depends on the number of factors $k$ and on $n$, where ($n\times n$) is the size of $\bm{A}$. We therefore seek to compute it for multiple values of $n$ ($n\in\{8000,12000\}$) and a range of values for $k$ (between 64 and 4000, with steps of 64), for each value of $n$. We use Theorem 6.2 to approximate it theoretically. To measure it empirically, we compute $\Delta_{0}^{prior}$ and $\hat{\Delta}_{0}$ based on $\bm{X}_{0}$ in the first step of the algorithm (it is independent of the matrix of interest $\bm{A}$), for $t$ runs of the algorithm. We can then use the $t$ measures to estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$. We claim that $t=5$ is enough to see a general tendancy, as we have estimates for many ($k,n$) couples. We can also compare the empirical estimates to the proposed asymptotic approximations specified in Approximation 6.2.
Figure 2 and Figure 3 (in Appendix) presents the comparison between empirical (blue dots) and theoretical (green line) estimates of $r(k,n)$ for $n=8000$ and $n=12000$. The empirical ratios are calculated from the averages over the $t=5$ runs for each value of $k$. The results indicate that both empirically and theoretically the proposed noise scaling $\hat{\Delta}_{l}$ is much tighter than $\Delta_{l}^{prior}$ at the first step of the algorithm. Our theoretical approximation is conservative and underestimates how much tighter the bound is initially, compared to what we observe in practice. The proposed bound is tighter by a multiplicative factor on the first step and can therefore drastically reduce the impact of the noise introduced by DP on the convergence of the Privacy Preserving power method, as the algorithm is usually run for very few steps ($L$ is usually in the range of $1$-$5$).
7 Conclusion and limitations
@@
We presented a differentially private and decentralized randomized power method that addresses privacy concerns in large-scale spectral analysis and recommendation systems. We introduced new convergence bounds, improving the accuracy of the method while ensuring privacy guarantees. By employing Secure Aggregation in a decentralized setting, we were able to reduce the noise introduced for privacy, maintaining the efficiency and privacy of the centralized version but adapting it for distributed environments.
@@
However, our privacy guarantees rely on the honest-but-curious threat model and assume no collusion or malicious behavior from participants, which may not always be realistic. In the future, it would also be interesting to analyze a private and decentralized version of the accelerated version of the randomized power method, which could potentially have a faster convergence.
8 Impact Statement (required by ICML)
Our contributions directly address the need for privacy-preserving algorithms in the context of large-scale machine learning systems, where sensitive personal or organizational data is often at stake. The decentralized implementation ensures that the method can be applied in some real-world environments where data is distributed across multiple users or devices, making it particularly suitable for recommender systems and other distributed applications. This work is expected to have implications for privacy-conscious machine learning in real-world systems, to perform large-scale data analysis while maintaining robust privacy protections.
@@ Appendix A Proofs of Results
We note $\bm{A}^{\prime}$ a matrix adjacent to $\bm{A}$ using (12).
@@
Although the proposed update model (12) is more general than when $\bm{A}^{\prime}$ is defined using (11), the proposed bound $\hat{\Delta}_{l}$ is also generally tighter than the bound proposed in (Hardt & Price, 2014; Balcan et al., 2016).
@@ A.2.1 Note on related privacy proofs:
Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by Hardt & Price (2014), Balcan et al. (2016), Hardt & Roth (2013), Hardt & Roth (2012). These papers mostly rely on the same proofs to establish that the PPM satisfies ($\epsilon$, $\delta$)-DP. Specifically, Balcan et al. (2016) references the privacy proof from Hardt & Price (2014), which builds upon the privacy proof from Hardt & Roth (2013). The proof therein is also closely related to the one of Hardt & Roth (2012).
However, both Hardt & Roth (2013) (Theorem 2.4) and Hardt & Roth (2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped. This error could potentially cause the privacy parameter $\epsilon$ in the proposed mechanism to be arbitrarily small, providing no privacy guarantee at all. The original composition rule is presented in Theorem III.3 of Dwork et al. (2010). Moreover, Lemma 3.4 of Hardt & Roth (2013) misuses Theorem 2.4. Indeed, they claim that their algorithm satisfies $(\epsilon^{\prime},\delta)$-DP at each iteration. By their Theorem 2.4, then the algorithm overall should satisfy $(\epsilon^{\prime},k\delta+\delta^{\prime})$-DP where $\delta^{\prime}>0$, and not $(\epsilon^{\prime},\delta)$-DP as claimed.
A.2.2 Proof of Theorem 3.2

### v3

@@ 5.1 Decentralized and private power method using distributed DP:
We introduce a federated version of the Privacy-Preserving Power Method in Algorithm 2. This version significantly reduces the noise variance by a factor of $sp\log n\,\|\bm{U}\|_{\infty}^{2}$ compared to the method in Balcan et al. (2016). The improvement comes from the fact that we emulate centralized Differential Privacy (DP) over secure channels, avoiding the higher noise required in local DP settings with public channels, and that we use our generally tighter sensitivity bound.
We now give a simplified definition of Secure Aggregation and demonstrate the equivalence between Algorithm 2 and our centralized Algorithm 1 in Theorem 5.2 (proof in Appendix A.4).
@@
Let $SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1{\leq}i{\leq}s\})$ the Secure Aggregation of matrices $\bm{Y}_{\ell}^{(i)}$ held by users indexed by $\{i|1{\leq}i{\leq}s\})$. It is equivalent to computing $\bm{Y}_{\ell}=\sum_{i=1}^{s}\bm{Y}_{\ell}^{(i)}$ over secure channels.
1:  Input: distributed matrices $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, private parameters $\epsilon,\delta$.
3:  Initialization: orthonormal $\bm{X}^{0}\in\mathbb{R}^{n\times p}$ by QR decomposition on a random Gaussian matrix $\bm{G}_{0}$; noise variance parameter $\nu=\epsilon^{-1}\sqrt{\frac{4L\log(1/\delta)}{s}}$;
4:  for $\ell=1$ to $L$ do
5:      1. The central node broadcasts $\bm{X}^{\ell-1}$ to all $s$ computing nodes;2. Computing node $i$ computes $\bm{Y}^{(i)}_{\ell}=\bm{A}^{(i)}\bm{X}^{\ell-1}+\bm{G}^{(i)}_{\ell}$ with $\bm{G}^{(i)}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\nu^{2})^{n\times p}$;3. The central node computes with the clients $\bm{Y}_{\ell}=SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1\leq i\leq s\})$. 4. The central node computes QR factorization $\bm{Y}_{\ell}=\bm{X}^{\ell}\bm{R}_{\ell}$.
6:  end for
@@ 5.2 Communication and computation cost analysis:
Table 1 presents the per-iteration communication and computation overhead introduced by both Secure Aggregation (Bell et al., 2020) and non-secure federated aggregation. We can see that Secure Aggregation introduces some non-dominant logarithmic terms.
@@ 6 Empirical comparison of the proposed bounds
We introduced runtime-dependent and runtime-independent convergence bounds for our algorithm. The runtime-dependent bound depends on $\bm{A}$ and the iteratively computed $\bm{X}^{l}$, whereas the runtime-independent bound (Theorem 4.1) only depends on $\bm{A}$. To illustrate the practical impact of our changes, we focus on an application in recommender systems and compare how our algorithms perform in this context. Additionally, we use a statistical approximation to see how the bounds compare at the first step of the algorithm (sketching step), regardless of the application in Appendix C.
@@ 6.1 Application to recommender systems
State-of-the-art recommender systems like GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) utilize singular or eigenvalue decomposition as part of their algorithms. Specifically, they represent the dataset (user-item interactions) using a normalized adjacency matrix $\tilde{\bm{R}}$, from which they compute a normalized item-item matrix $\tilde{\bm{P}}$. To get rid of noise in $\tilde{\bm{R}}$, they compute an ideal low-pass filter based on the top-$p$ eigenvectors of $\tilde{\bm{P}}$ and apply it to $\tilde{\bm{R}}$. Detailed definitions are provided in Appendix B.
The first four columns of Table 2 presents key statistics of popular recommendation datasets used in the literature. Depending on the use case, these datasets may be held by a central curator, or distributed among the users of a recommender system. Accordingly, either the centralized PPM (Algorithm 1) or the decentralized PPM (Algorithm 2) can be applied to compute the leading eigenvectors. To demonstrate the practicality of our proposed methods, we focus on a decentralized setting where each user has access to its own ratings and collaborates with other users to compute the desired eigenvectors. $\Delta_{l}^{prior}=\mu_{0}(\bm{A})\sqrt{p\cdot\log({n})}$ denotes the sensitivity bound proposed in (Balcan et al., 2016) while ${\hat{\Delta}_{l}}=\mu_{1}(\bm{A})$ uses our multidimensional refinement.
@@
Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{prior}$ (prior) ${\hat{\Delta}_{l}}$ (ours)
Amazon-book 52,643 91,599 2,984,108 0.33 0.99 1.12 ×$\times\sqrt{p}$ 0.99
MovieLens 71,567 10,677 7,972,582 0.39 1.00 1.19 ×$\times\sqrt{p}$ 1.00
EachMovie 74,425 1,649 2,216,887 0.49 1.00 1.33 ×$\times\sqrt{p}$ 1.00
Jester 54,906 151 1,450,010 0.73 1.00 1.64 ×$\times\sqrt{p}$ 1.00
(a) MovieLens dataset.
(b) EachMovie dataset.
Figure 1: Comparison of the impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}},$ where $\tilde{\bm{R}_{p}}$ represents the approximated matrix and $\bm{R}_{p}$ the original matrix. Results are means on 10 runs with shaded bands indicating 99% confidence intervals (computed via bootstrap), shown for the MovieLens and EachMovie datasets, with $p=32$ and $L=3$.
Empirical comparison of the runtime-independent bounds.
As discussed in Section 4, our proposed runtime-independent bound is tighter than those in Hardt and Price (2014); Balcan et al. (2016) when $\mu_{1}(\bm{A})\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({n})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$. The last two columns from Table 2 show that this condition holds for popular recommender system datasets. As we can deduce from Table 2, the proposed bound theoretically allows us to converge to solutions with much smaller $\eta$ values, especially when the desired number of factors $p$ is large.
Practical utility of the proposed algorithms.
We saw that the proposed runtime-independent bounds were practically tighter than the previous one for the task of interest. We now demonstrate that Algorithm 2 can be used to compute the top-$p$ eigenvectors $\bm{U}_{p}$ of the item-item matrix $\tilde{\bm{P}}$ under ($\epsilon$,$\delta$)-DP (proof in Appendix B.1):
@@
Algorithm 2 with $\Delta_{l}=\sqrt{2}\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ can approximate $\bm{U}_{p}$, the top-$p$ eigenvectors of $\tilde{\bm{P}}$ in a decentralized setting under $(\epsilon,\delta)$-Differential Privacy.
@@
Both GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) use $\bm{U}_{p}$ to compute the ideal low pass filter and apply it to the interaction matrix $\bm{R}$, yielding $\bm{R}_{p}$. To illustrate the practicality of our proposed method, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using either our decentralized PPM, the previous PPM versions in  Balcan et al. (2016) or FedPower from Guo et al. (2024). We then use $\tilde{\bm{U}}_{p}$ to compute an approximation of $\bm{R}_{p}$ (denoted by $\tilde{\bm{R}_{p}}$), and compute the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to each version. We provide experimental details in Appendix B.2.
Figure 1 illustrates the impact of the privacy parameter $\epsilon$ on the approximation error for the EachMovie and MovieLens datasets, with $p=32$ and $L=3$, clearly showing the advantage of our method over prior work. Our decentralized PPM achieves relative approximation error of $\approx 1/10$ for values of $\epsilon\in(5,10)$ for EachMovie and $\epsilon\approx 20$ for MovieLens. In contrast, other decentralized methods (Balcan et al., 2016; Guo et al., 2024) require $\epsilon$ to be of the order of at least $10^{3}$ to achieve comparable errors, which does not seem to provide meaningful privacy protection. [^3] We refer to Dwork et al. (2019) for more practical details on how to set $\epsilon$ and provide additional comparisons and details in Appendix B.1. (Guo et al., 2024) yields the worst relative approximation errors. We hypothesize that this is because it uses worst case, non-adaptive sensitivity bounds for DP, as opposed to our proposed method and those of Balcan et al. (2016). For a fixed approximation error, both our centralized and decentralized methods yield $\epsilon$ roughly four times smaller than required by the centralized method of Balcan et al. (2016), demonstrating that our propositions significantly strengthen privacy guarantees.
7 Conclusion
@@
We presented differentially private versions of the centralized and decentralized randomized power method that addresses privacy concerns in large-scale spectral analysis and recommendation systems. We introduced a new sensitivity bound, which we show theoretically and empirically to improve the accuracy of the method while ensuring privacy guarantees. By employing Secure Aggregation in a decentralized setting, we can reduce the noise introduced for privacy, maintaining the efficiency and privacy of the centralized version but adapting it for distributed environments. Our methods could enable organizations (from healthcare networks to social platforms) to extract useful structure from data without compromising individual records.
@@
Limitations. Our privacy guarantees require no dropout from participants to ensure distributed DP, which may not always be realistic. It would also be interesting to analyze a private and decentralized version of the accelerated version of the randomized power method, which could potentially converge faster.
@@ Appendix A Proofs of Results
We denote by $\bm{A}^{\prime}$ a matrix adjacent to $\bm{A}$ using (7):
@@
Although the proposed update model (7) is more general than when $\bm{A}^{\prime}$ is defined using (6), the proposed bound $\hat{\Delta}_{l}$ is also generally tighter than the bound proposed in Hardt and Price (2014); Balcan et al. (2016).
@@ A.2.1 Note on related privacy proofs:
Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by Hardt and Roth (2012, 2013); Hardt and Price (2014); Balcan et al. (2016). These papers mostly rely on the same proofs to establish that the PPM satisfies ($\epsilon$, $\delta$)-DP. Specifically, Balcan et al. (2016) references the privacy proof from Hardt and Price (2014), which builds upon the privacy proof from Hardt and Roth (2013). The proof therein is also closely related to the one of Hardt and Roth (2012).
However, both Hardt and Roth (2013) (Theorem 2.4) and Hardt and Roth (2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped. This error could potentially cause the privacy parameter $\epsilon$ in the proposed mechanism to be arbitrarily small, providing no privacy guarantee at all. The original composition rule is presented in Theorem III.3 of Dwork et al. (2010). Moreover, Lemma 3.4 of Hardt and Roth (2013) misuses Theorem 2.4. Indeed, they claim that their algorithm satisfies $(\epsilon^{\prime},\delta)$-DP at each iteration. By their Theorem 2.4, then the algorithm overall should satisfy $(\epsilon^{\prime},k\delta+\delta^{\prime})$-DP where $\delta^{\prime}>0$, and not $(\epsilon^{\prime},\delta)$-DP as claimed.
A.2.2 Zero-Concentrated Differential Privacy (zCDP):
$D_{\alpha}(\mathcal{M}(D_{1})\|\mathcal{M}(D_{2}))\leq\rho\alpha\,,$ (17)
The following lemma, introduced by Bun and Steinke (2016), specifies how the addition of Gaussian noise can be used to construct a randomized algorithm that satisfies zCDP.
Lemma A.1 (Gaussian Mechanism (Proposition 1.6 (Bun and Steinke, 2016))).
The next lemma, which is a generalized version of a result presented by Bun and Steinke (2016), explains how a randomized algorithm, constructed by recursively composing a sequence of zCDP-satisying randomized algorithms, also satisfies zCDP.
Lemma A.2 (Adaptive composition (Generalization from Lemma 2.3 of (Bun and Steinke, 2016))).
Then $M^{\prime\prime}$ satisfies $(\sum_{i=1}^{L}\rho_{i})$-zCDP.
A.2.3 Proof of Theorem 3.2